OpenStudy (anonymous):

f(x)=3(x-3)^2 The vertex is_____ The equation of the of symmetry is_____

OpenStudy (anonymous):

\[(y-0)= 4\times \frac 34 (x-3)^2 \] Vertex is at \((3,0)\) Equation of the line of symmetry is \(x=3\)

OpenStudy (anonymous):

Wow I was way off on this one...thanks!!